# Vortices and Factorization

### v2

@@ 3 $\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
where we used (17). In difference from the Adler-Moser case, there are two distinct branches of polynomials both starting at $p_{0}=1$ and $q_{0}=1$. One branch with $i\geq 0$ going in the positive direction
$\begin{array}[]{ll}q_{0}=1&\quad p_{0}=1\\ q_{1}=z&\quad p_{1}=z^{5}+t_{1}\\ q_{2}=z^{5}+s_{2}z-4t_{1}&\begin{array}[]{r}\;\;p_{2}={z}^{16}+{\frac{44}{7}}s_{{2}}{z}^{12}-32t_{{1}}{z}^{11}+22{s_{{2}}}^{2}{z}^{8}-{\frac{2112}{7}}t_{{1}}s_{{2}}{z}^{7}\\ +1408{t_{{1}}}^{2}{z}^{6}+t_{{2}}{z}^{5}-44{s_{{2}}}^{3}{z}^{4}+352t_{{1}}{s_{{2}}}^{2}{z}^{3}\\ -1408s_{{2}}{t_{{1}}}^{2}{z}^{2}+2816{t_{{1}}}^{3}z+t_{{2}}t_{{1}}-{\frac{11}{5}}{s_{{2}}}^{4}\end{array}\\ \dots&\dots\end{array}$ (25)
etc, and another branch with $i\leq 0$ going in the negative direction
$\begin{array}[]{lll}p_{0}=1&&q_{0}=1\\ p_{-1}=z&&q_{-1}=z^{2}+s_{-1}\\ \begin{aligned} p_{-2}=z^{8}+\frac{28}{5}s_{-1}z^{6}+14s_{-1}^{2}z^{4}\\ +28s_{-1}^{3}z^{2}+t_{-2}z-7s_{-1}^{4}\end{aligned}&&\begin{aligned} q_{-2}=z^{7}+7s_{-1}z^{5}+35s_{-1}^{2}z^{3}+s_{-2}z^{2}\\ -35s_{-1}^{3}z+s_{-1}s_{-2}-\frac{5}{2}t_{-2}\end{aligned}\\ \dots&&\dots\end{array}$ (26)
etc, where $t_{i}$ and $s_{i}$ stand for arbitrary complex constants  [^2] Without loss of generality we omit a constant corresponding to a shift of $z$, so that $q_{1}=p_{-1}=z$. emerging in course of integrations in (18,19) or (20).
Summarizing, we emphasise again that, for $\Lambda=1$ and $\Lambda=2$, any no-common root solutions of the bi-linear equation (8) can be obtained by the recursive procedure described in this section. Therefore, the main sequences (23) or (26,25) determine all possible static patterns of a finite number of vortices with ${\cal Q}_{i}\in\{-1,1\}$ or ${\cal Q}_{i}\in\{-1,2\}$ respectively.
@@ 4 Darboux Transformations and Factorization Chains
The above results can be alternatively obtained with the help of factorization methods. The approach relying on these methods is more general than that used in the previous section and allows to obtain new stationary configurations. It involves theory of one-dimensional Schrodinger operator (“quantum Hamiltonian”) with potential $u(z)$:
@@
Let us now take the operator
$H-\eta=-\partial_{x}^{2}+u(z)-\eta,$ (27)
and factorize into the product of two first-order operators
$H-\eta=A^{*}A,$ (28)
where [^3] Note that we do not use the special symbol $\circ$ for composition of operators and instead e.g. $A^{*}\circ A$ or $\varkappa\circ\partial_{z}\circ\varkappa^{-1}$ etc we simply write $A^{*}A$ or $\varkappa\partial_{z}\varkappa^{-1}$ etc.
@@
The function $\varkappa=\varkappa(z)$ is an arbitrary element of kernel of $H-\eta$, i.e. $\varkappa$ is any eigenfunction of $H$ corresponding to the eigenvalue $\eta$
@@
It trivially follows from (28) that for any $\lambda$
$A(H-\lambda)=(\hat{H}-\lambda)A,$ (30)
where the new Schrodinger operator $\hat{H}$ is obtained from $H$ by permutation of factors $A^{*}$ and $A$ in (28):
$\hat{H}=AA^{*}+\eta=-\partial_{x}^{2}+\hat{u},\quad\hat{u}=u-2(\log\varkappa)^{\prime\prime}.$ (31)
Now, let us take an eigenfunction $\psi$ corresponding to the eigenvalue $\lambda$ of the old Schrodinger operator, i.e. $H\psi=\lambda\psi$. Then, from (30) it easy to see that the following transform of $\psi$:
@@
is an eigenfunction of the new operator $\hat{H}$ corresponding to the same eigenvalue $\lambda$, i.e.
@@
In other words, we can obtain a new operator $\hat{H}$ and its eigenfunctions $\hat{\psi}(z;\lambda)$ from the old operator $H$ and its eigenfunctions $\psi(z;\lambda)$ corresponding to the same eigenvalues $\lambda$. New operator and eigenfunctions are obtained with the help of an eigenfunction $\varkappa=\psi(z;\eta)$ of the old operator $H$. The transformation $H\to\hat{H}$, $\psi\to\hat{\psi}$ given by (31,32) is called the Darboux Transformation.
When $\lambda\not=\eta$ the transform (32) of the two dimensional kernel of $H-\lambda$ is also two-dimensional. However, when $\lambda=\eta$, the transformation (32) annihilates the eigenfunction $\varkappa$ of $H$, i.e. $A\varkappa=0$: It maps the two-dimensional kernel $C_{1}\varkappa+C_{2}\varkappa\int\frac{dz}{\varkappa^{2}}$ of $H-\eta$ to the one dimensional sub-space
@@
where the primitive of $\varkappa^{2}$ includes an arbitrary constant of integration. This extends the Darboux transformation of an eigenfunction (32) to the case $\lambda=\eta$. Transformations (33) are usually considered for $\eta=0$ and are called the Darboux transformations at the zero energy level.
Darboux transformations can be iterated. Sequence of iterated transformations is presented by the factorization chain:
$H_{0}=A_{0}^{*}A_{0}+\eta_{1}\to H_{1}=A_{0}A_{0}^{*}+\eta_{1}=A_{1}^{*}A_{1}+\eta_{2}\to H_{2}=A_{1}A_{1}^{*}+\eta_{2}=A_{2}^{*}A_{2}+\eta_{3}\to\dots.$ (34)
When all $\eta_{i}=0$, (34) becomes the chain of transformations at the zero energy level [^4] It would be more correctly to call it the chain of the Darboux transformations associated with zero energy level, rather than at the zero energy level, since eigenfunctions at all energy levels are transformed: Those with zero eigenvalue transform according to (33) and the rest according to (32).
@@
According to (33) and (31)
$\varkappa_{i+1}=\frac{C_{i}}{\varkappa_{i}}\int\varkappa_{i}^{2}dz,,\quad u_{i+1}=u_{i}-2(\log\varkappa_{i})^{\prime\prime},\quad H_{i}\varkappa_{i}=0,$ (36)
where $C_{i}$ are arbitrary constants. Using (36), we can construct iteratively the chain (35). Note, that at each step of the chain a free parameter, which is a constant of integration, appears in (36).
It trivially follows from (34) that $H_{0}$ and $H_{n}$ are related by the intertwining operator $T_{n}$,
@@
The function $T_{n}\psi(x;\lambda)$, where $\psi$ is an eigenfunction of $H_{0}$, ia an eigenfunction of $H_{n}$. Due to the Crum therem [10], the intertwining operator as well as potential $u_{n}$ can be explicitly written down in terms of eigenfunctions of the initial hamiltonian $H_{0}$. In more detail, the theorem states that the composition of $n$ Darboux transformations, corresponding to the factorization chain (34), transforms the eigenfunction $\psi$, corresponding to the eigenvalue $\lambda$ of $H_{0}=-\partial_{x}^{2}+u_{0}$ to the eigenfunction $\hat{\psi}$ of $H_{n}=-\partial_{x}^{2}+u_{n}$, corresponding to the same eigenvalue, as the ratio of two wronskians
$\hat{\psi}=\frac{{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n},\psi]}{{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n}]}$ (38)
while the potential is transformed in the following way
$u_{n}=u_{0}-2\left(\log{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n}]\right)^{\prime\prime}.$ (39)
Here $\psi_{i}$ are $n$ eigenfunctions of the initial Schrodinger operator $H_{0}$ corresponding to the distinct eigenvalues $\eta_{1},\eta_{2},\dots,\eta_{n}$, i.e. $H_{0}\psi_{i}=\eta_{i}\psi_{i}$, $1\leq i\leq n$, called “seed functions”, and ${\cal W}$ stands for the Wronskian determinant. From (38), we get the Wronskian representation of the intertwining operator [^5] Simplest proof of the Crum theorem uses the fact that, according to (32,37), $T_{n}$ annihilates $n$ seed functions $\psi_{i}$, $i=1,\dots,n$, i.e. kernel of $T_{n}$ is spanned by these seed functions. Then, since $T_{n}=\partial_{z}^{n}+\dots$, we immediately come to (40).
$T_{n}[\,\boldsymbol{\cdot}\,]=\frac{{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n},\,\boldsymbol{\cdot}\,]}{{\cal W}[\psi_{1},\psi_{2},\dots,\psi_{n}]}.$ (40)
In the particular case of confluent eigenvalues, when the spacing between $\eta_{i}$ vanishes, i.e. all $\eta_{i}$ tend to one value $\eta$ and we have the chain (35), the Crum theorem degenerates to
@@
where $\varkappa=\varkappa(z;\eta)$ is an eigenfunction of $H_{0}$ corresponding to the eigenvalue $\eta$, and $\varkappa_{n}$ is an eigenfunction of $H_{n}$ with the same eigenvalue. Here, the intertwining operator degenerates to
@@
Eq. (41) is useful for representing the Adler-Moser polynomials in the Wronskian form. Indeed, let us take derivatives of the Schrodinger equation $H_{0}\varkappa(z,\eta)=\eta\varkappa(z,\eta)$, $H_{0}=-\partial_{z}^{2}+u_{0}(z)$ wrt $\eta$ and then set $\eta=0$. In this way, we get the recurrence relation for derivatives $\partial^{n}_{\eta}\varkappa$ at zero energy level:
@@
$\varkappa_{n}=\frac{\tau_{n}}{\tau_{n-1}},\quad\tau_{n}={\cal W}[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{n}],\quad u_{n}=u_{0}-2(\log\tau_{n-1})^{\prime\prime},\quad\tau_{0}=1,$ (42)
@@
$H_{0}{\cal X}_{n}=c_{n}{\cal X}_{n-1},\quad H_{0}{\cal X}_{1}=0,\quad H_{0}=-\partial_{z}^{2}+u_{0}(z).$ (43)
In eq. (43), $c_{n}$ are arbitrary constants. At each iteration in (43) two integration constants appear, but only one of them is essential. The intertwining operator writes in terms of ${\cal X}_{i}$ as
@@
Here [^6] In (45) we set $c_{n}=2n+1$. For such choice of normalization constants, $\tau_{n}$ is monic in $z$., the function ${\cal X}_{n}$ is a polynomial of degree $12n-$ 1 in $z$. Therefore $\tau_{n}$ in (42) is a polynomial of degree $n(n+1)/2$. Substituting $\varkappa_{n}=\tau_{n}/\tau_{n-1}$ (see(42)) into (36), we obtain recurrence relation for $\tau_{n}$ which coincides with that (22) for the Adler-Moser polynomials $P_{n}$. Taking into account that degrees of $\tau_{n}$ and $P_{n}$ are the same, we come to the conclusion that $\tau_{n}=P_{n}$. Thus, we get the Wronskian representation for the Adler-Moser polynomials

### v3

@@ 3 $\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
where we used (18). Unlike the Adler-Moser case, there are two distinct branches of polynomials, both starting at $p_{0}=1$ and $q_{0}=1$. One branch, with $i\geq 0$, goes in the positive direction:
$\begin{array}[]{ll}q_{0}=1&\quad p_{0}=1\\ q_{1}=z&\quad p_{1}=z^{5}+r_{1}\\ q_{2}=z^{5}+s_{2}z-4r_{1}&\begin{array}[]{r}\;\;p_{2}={z}^{16}+{\frac{44}{7}}s_{{2}}{z}^{12}-32r_{{1}}{z}^{11}+22{s_{{2}}}^{2}{z}^{8}-{\frac{2112}{7}}r_{{1}}s_{{2}}{z}^{7}\\ +1408{r_{{1}}}^{2}{z}^{6}+r_{{2}}{z}^{5}-44{s_{{2}}}^{3}{z}^{4}+352r_{{1}}{s_{{2}}}^{2}{z}^{3}\\ -1408s_{{2}}{r_{{1}}}^{2}{z}^{2}+2816{r_{{1}}}^{3}z+r_{{2}}r_{{1}}-{\frac{11}{5}}{s_{{2}}}^{4}\end{array}\\ \dots&\dots\\ q_{n}=q_{n}(z;r_{1},r_{2},\dots r_{n-1};s_{2},s_{3},\dots s_{n})&\quad p_{n}=p_{n}(z;r_{1},r_{2},\dots,r_{n};s_{2},s_{3},\dots s_{n})\\ \dots&\dots\end{array}$ (26)
etc, and another branch, with $i\leq 0$, goes in the negative direction:
$\begin{array}[]{ll}p_{0}=1&\quad q_{0}=1\\ p_{-1}=z&\quad q_{-1}=z^{2}+s_{-1}\\ \begin{array}[]{r}p_{-2}=z^{8}+\frac{28}{5}s_{-1}z^{6}+14s_{-1}^{2}z^{4}\\ +28s_{-1}^{3}z^{2}+r_{-2}z-7s_{-1}^{4}\end{array}&\quad\begin{array}[]{r}q_{-2}=z^{7}+7s_{-1}z^{5}+35s_{-1}^{2}z^{3}+s_{-2}z^{2}\\ -35s_{-1}^{3}z+s_{-1}s_{-2}-\frac{5}{2}r_{-2}\end{array}\\ \dots&\dots\\ p_{-n}=p_{-n}(z;r_{-2},r_{-3}\dots r_{-n};s_{-1},s_{-2},\dots s_{1-n})&\quad q_{-n}=q_{-n}(z;r_{-2},r_{-3}\dots r_{-n};s_{-1},s_{-2},\dots s_{-n})\\ \dots&\dots\end{array}$ (27)
etc, where $r_{i}$ and $s_{i}$ are arbitrary complex constants  [^7] Without loss of generality we omit a constant corresponding to a shift of $z$, so that $q_{1}=p_{-1}=z$. emerging in the course of integrations in (19,20) or (21). Here, we use the following normalization for the constants:
$q_{\pm n}=\tilde{q}_{\pm n}+s_{\pm n}q_{\pm(n-1)},\quad p_{\pm n}=\tilde{p}_{\pm n}+r_{\pm n}p_{\pm(n-1)},$ (28)
where the term proportional to $z^{\deg q_{\pm(n-1)}}$ is absent in $\tilde{q}_{\pm n}$ and the term proportional to $z^{\deg p_{\pm(n-1)}}$ is absent in $\tilde{p}_{\pm n}$.
Summarizing, we emphasize again that for $\Lambda=1$ and $\Lambda=2$, any solutions of the bi-linear equation (10) with no common/multiple roots can be obtained through the recursive procedure described in this section. Therefore, the main sequences (24) or (27,26) determine all possible static patterns of a finite number of vortices with ${\cal Q}_{i}\in\{-1,1\}$ or ${\cal Q}_{i}\in\{-1,2\}$, respectively.
@@ 4 Darboux Transformations and Factorization Chains
The results above can also be obtained using factorization method. This approach not only recovers the previous results, but also enables the construction of new families of stationary configurations. In this section, we review the factorization method for Schrodinger operators. Third-order operators will be considered in Sections 8 and 9.
The Schrodinger operator (“quantum Hamiltonian”) with potential $u(z)$ has the form:
@@
We can express it as
$H=A^{*}A+\eta,$ (29)
where $\eta$ is a constant and $A$, $A^{*}$ are formally adjoint first-order operators [^8] Note that we do not use the special symbol $\circ$ for the composition of operators. Instead, we write expressions like $A^{*}A$ or $\varkappa\partial_{z}\varkappa^{-1}$, without using the composition symbol.:
@@
The function $\varkappa=\varkappa(z)$ is an eigenfunction of $H$ corresponding to the eigenvalue $\eta$
@@
By permuting factors $A^{*}$ and $A$ in (29), we get the new Schrodinger operator $\hat{H}$
$\hat{H}=AA^{*}+\eta=-\partial_{z}^{2}+\hat{u},\quad\hat{u}=u-2(\log\varkappa)^{\prime\prime}.$ (31)
From (29) and (31) it follows that for any constant $\lambda$
$A(H-\lambda)=(\hat{H}-\lambda)A.$ (32)
Now, let us take an eigenfunction $\psi$ corresponding to the eigenvalue $\lambda$ of the original Schrodinger operator, i.e. $H\psi=\lambda\psi$. Then, from (32) we see that following transform of $\psi$:
@@
is an eigenfunction of the new operator $\hat{H}$ corresponding to the same eigenvalue $\lambda$, i.e.,
@@
In other words, we can obtain a new operator $\hat{H}$ and its eigenfunctions $\hat{\psi}=\hat{\psi}(z;\lambda)$ from the old operator $H$ and its eigenfunctions $\psi=\psi(z;\lambda)$ corresponding to the same eigenvalues $\lambda$. New operator and eigenfunctions are obtained with the help of a “seed” eigenfunction $\varkappa=\psi(z;\eta)$ of the old operator $H$, corresponding to the “seed” eigenvalue $\eta$. The transformation $H\to\hat{H}$, $\psi\to\hat{\psi}$ given by (31,33) is called the Darboux Transformation.
For all eigenvalues $\lambda$, except the seed eigenvalue, i.e. for $\lambda\not=\eta$, the transform (33) of the two dimensional eigenspace of $H$ is also two-dimensional. However, when $\lambda=\eta$, the transformation (33) annihilates the eigenfunction $\varkappa=\psi(z;\eta)$ of $H$, i.e. $A\varkappa=0$: It maps the two-dimensional kernel $C_{1}\varkappa+C_{2}\varkappa\int\frac{dz}{\varkappa^{2}}$ of $H-\eta$ to the one dimensional sub-space
@@
where $C$ is an arbitrary constant and the primitive of $\varkappa^{2}$ includes another arbitrary constant of integration. This extends transformation of an eigenfunction (33) to the case $\lambda=\eta$. Iterations of transformation (34) are typically considered for $\eta=0$. In such cases, transformations (34) are reffered to as Darboux transformations at the zero energy level.
In general, the sequence of iterated transformations is presented by the factorization chain, usually referred to as the chain of Darboux transformations at levels  [^9] It would be more accurate to refer to it as the chain of Darboux transformations associated with levels $\lambda=\eta_{i}$, rather than at levels $\lambda=\eta_{i}$, since eigenfunctions at all values of $\lambda$ are transformed. $\lambda=\eta_{i}$:
$H_{0}=A_{0}^{*}A_{0}+\eta_{0}\to H_{1}=A_{0}A_{0}^{*}+\eta_{0}=A_{1}^{*}A_{1}+\eta_{1}\to H_{2}=A_{1}A_{1}^{*}+\eta_{1}=A_{2}^{*}A_{2}+\eta_{2}\to\dots.$ (35)
When all $\eta_{i}=0$, (35) becomes the chain of transformations at the zero energy level:
@@
According to (34) and (31),
[…]
where $C_{i}$ are arbitrary multiplication constants, and the primitives include arbitrary constants of integration. Using (37), we can construct iteratively the chain (36). At each step of the chain a free parameter (arising from the constant of integration) appears in (37).
It follows from (35) that $H_{0}$ and $H_{n}$ are related by the intertwining operator [^10] Additionally, conjugate (“inverse”) intertwining identities hold: $T^{*}_{n}H_{n}=H_{0}T^{*}_{n}$, $T^{*}_{n}=A^{*}_{0}\cdots A^{*}_{n-2}A^{*}_{n-1}$. $T_{n}$:
@@
The function $T_{n}\psi(x;\lambda)$, where $\psi$ is an eigenfunction of $H_{0}$, is an eigenfunction of $H_{n}$. According to the Crum therem [15], the intertwining operator as well as potential $u_{n}$ can be explicitly written down in terms of eigenfunctions of the initial hamiltonian $H_{0}$. In more detail, the theorem states that the composition of $n$ Darboux transformations, corresponding to the factorization chain (35), transforms the eigenfunction $\psi$, corresponding to the eigenvalue $\lambda$ of $H_{0}=-\partial_{z}^{2}+u_{0}$ to the eigenfunction $\hat{\psi}$ of $H_{n}=-\partial_{z}^{2}+u_{n}$, corresponding to the same eigenvalue, as the ratio of two Wronskians
$\hat{\psi}=\frac{{\cal W}[\psi_{0},\psi_{1},\dots,\psi_{n-1},\psi]}{{\cal W}[\psi_{0},\psi_{1},\dots,\psi_{n-1}]}.$ (39)
Meanwhile, the potential is transformed in the following way
$u_{n}=u_{0}-2\left(\log{\cal W}[\psi_{0},\psi_{1},\dots,\psi_{n-1}]\right)^{\prime\prime}.$ (40)
Here $\psi_{i}$ are $n$ eigenfunctions of the initial Schrodinger operator $H_{0}$ corresponding to the distinct eigenvalues $\eta_{0},\eta_{2},\dots,\eta_{n-1}$, i.e. $H_{0}\psi_{i}=\eta_{i}\psi_{i}$, $10\leq i\leq n-$ 1, called “seed functions”, and ${\cal W}$ stands for the Wronskian determinant. From (39), we get the Wronskian representation of the intertwining operator [^11] Simplest proof of the Crum theorem uses the fact that, according to (33,38), $T_{n}$ annihilates $n$ seed functions $\psi_{i}$, $i=0,\dots,n-1$, i.e. kernel of $T_{n}$ is spanned by these seed functions. Then, since $T_{n}=\partial_{z}^{n}+\dots$, we immediately come to (41).
$T_{n}[\,\boldsymbol{\cdot}\,]=\frac{{\cal W}[\psi_{0},\psi_{1},\dots,\psi_{n-1},\,\boldsymbol{\cdot}\,]}{{\cal W}[\psi_{0},\psi_{1},\dots,\psi_{n-1}]}.$ (41)
In the particular case of confluent eigenvalues, when the spacing between $\eta_{i}$ vanishes, all $\eta_{i}$ tend to the same value $\eta$, and we have the chain of the type (36), the Crum theorem degenerates to:
@@
where $\varkappa=\varkappa(z;\eta)$ is an eigenfunction of $H_{0}$ corresponding to the eigenvalue $\eta$, and $\varkappa_{n}$ is an eigenfunction of $H_{n}$ with the same eigenvalue. Here, the intertwining operator degenerates to :
@@
Eq. (42) is useful for representing the Adler-Moser polynomials in the Wronskian form. Indeed, let us take derivatives of the Schrodinger equation $H_{0}\varkappa(z,\eta)=\eta\varkappa(z,\eta)$, $H_{0}=-\partial_{z}^{2}+u_{0}(z)$ wrt $\eta$ and then set $\eta=0$. In this way, we obtain the recurrence relation for derivatives $\partial^{n}_{\eta}\varkappa$ at the zero energy level:
@@
$\varkappa_{n}=\frac{\tau_{n+1}}{\tau_{n}},\quad\tau_{n}={\cal W}[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{n-1}],\quad u_{n}=u_{0}-2(\log\tau_{n})^{\prime\prime},\quad\tau_{0}=1,$ (43)
@@
$H_{0}{\cal X}_{n}=c_{n}{\cal X}_{n-1},\quad H_{0}{\cal X}_{1}=0,\quad H_{0}=-\partial_{z}^{2}+u_{0}(z),$ (44)
In eq. (44), $c_{n}$ are arbitrary constants. At each iteration of the Darboux transformation (37), apart from the common factor $C_{i}$, only one integration constant appear. Therefore, although, apart from $c_{n}$, two integration constant appear at each iteration of (44), only one of them is esential [^12] In other words, the sequence ${\cal X}_{1}$, ${\cal X}_{2}+c_{2}{\cal C}_{1}{\cal X}_{1}$, ${\cal X}_{3}+c_{3}{\cal C}_{1}{\cal X}_{2}+c_{3}c_{2}{\cal C}_{2}{\cal X}_{1},\dots$, where ${\cal C}_{i}$ are arbitrary constants, is also a solution of (44), and the values of Wronskians of this sequence ${\cal W}[{\cal X}_{1},{\cal X}_{2}+c_{2}{\cal C}_{1}{\cal X}_{1},\dots]$ do not depend on ${\cal C}_{i}$..
The intertwining operator is expressed in terms of ${\cal X}_{i}$ as
@@
Here [^13] In (46) we set $c_{n}=2n+1$. With this choice of normalization constants, $\tau_{n}$ is monic in $z$., the function ${\cal X}_{n}$ is a polynomial of degree $12n-$ 1 in $z$. Substituting $\varkappa_{n}=\tau_{n}/\tau_{n-1}$ (see(43)) into (37), we obtain recurrence relation for $\tau_{n}$ which coincides with that (23) for the Adler-Moser polynomials $P_{n}$. Since the initial conditions of recurrence $\tau_{0}=P_{0}=1$, $\tau_{1}=P_{1}=z$ are the same for $\tau_{n}$ and $P_{n}$, we come to the conclusion that $\tau_{n}=P_{n}$. Thus, we obtain the Wronskian representation for the Adler-Moser polynomials